Zoom Meeting

CRY+VIS Interface Meeting

Date: 2021/8/20 13:00-13:30

Participants: Takahashi, Node, Ikeda, Ushiba, Aso, Sato, Uchiyama

Post O3 Upgrade portal




VIS Meeting




This is generated by python script in /users/Miyo/dropbox/Script/scrapping/

Summary Report





  1. Fabián set up the optical lever for SR3 and calibrated the QPDs

  2. I implemented the calibration factors and I derived the sensing matrix OPLEV2EUL using old data and diagonalized the optical levers using the SENSALIGN matrix.

  3. I reset the sensing matrix LVDT2EUL and actuation matrix COIL2EUL at the IP using Marks's old derivation because I changed it two years ago and I don't think they are reliable. Note that I didn't do diagonalization for these.

  4. I measured and fit transfer functions for all 15 degrees of freedom. (klog pending). All models were updated in the control filter banks as FM10.
  5. I made control filters base on some simple criteria (IP: DCCTRL, GAS: DAMP, IM: DAMP, TM: OLDAMP). (klog pending)
    • Filter bank convention: suspension model in FM10, -1 in FM9, DAMP filter in FM4, and if applicable, DC filters in FM3, and notches at FM5.
    • We don't have parallel control banks to separate DC and damping control. These DC filters work together with the DAMP filter and are not expected to work alone although it's stable to the DC filters turned out without the DAMP filters. DAMP filters, on the other hand, can work alone for damping.
  6. As of now, all loops are closed for 4 hours+ and work fine without issues. Furthermore, it survived the Vanuatu 6.8 magnitude earthquake a few hours ago (all seismic monitors went red at KAGRA).

Something to note 1. LVDTs of the GAS filters are very noisy especially at above 10 Hz. I see strange responses above 10 Hz. I have made additional low-pass filters for these controls.

  1. There's no windshield for the optical lever. So, the optical lever readouts, especially longitudinal, can be very noisy.
  2. I observed that the yaw chain mode was not very well damped. This mode is supposed to be damped by the passive damper. We can see very slow yaw oscillation right after the IP yaw controls are engaged.
  3. There's no inertial damping. I didn't touch the IDAMP filter bank and the BLEND filter bank at the IP.
  4. There's no IM2TM DC control. I can implement this if this is requested. However, note that there's no dedicated setpoint for this. Alternatively, the OPTICALIGN block can be used to offset the IM and hence the TM for alignment.
  5. Guardian has changed a lot and it's not maintained by me anymore. I don't how what's the current situation so I did not test it.
  6. Everything is reproducible from my jupyter notebook at /kagra/Dropbox/Subsystems/VIS/vis_commissioning/sr3/notebook.





See Schedule


2021/03/29 updated



Travel Plans

The week from 8/23

The week from 8/30

Next meeting

KAGRA/Subgroups/VIS/VISMinutes20210820 (last edited 2021-08-20 16:39:15 by RyutaroTakahashi)